OpenStudy (fanduekisses):

Converting 8.3 in^2 to mm^2 check my work please?

OpenStudy (fanduekisses):

\[8.3 inches.^{2} x \frac{ 25.4 mm }{ 1 inch }\] squared

OpenStudy (fanduekisses):

I get 5354.83 mm^2

OpenStudy (danjs):

all i can remember from inch to metric is 1 in = 2.54 cm, notice how to deal with the square terms, use the linear relationship 1 in = 2.54 cm, then square both sides ( 1 in ) ^2 = (2.54 cm) ^2 1 in^2 = 2.54^2 cm^2 

OpenStudy (danjs):

Then if you can do it in your head, mm to cm is 10^-1, or you can always just convert to base units meters first, so all you need to remember is how the base unit relates to each one 1 cm = 10^-2m 1 mm = 10^-3m

OpenStudy (danjs):



OpenStudy (danjs):

remember to square both the conversion number and the unit [ 2.54 cm ] ^2 = 6.45 cm^2

OpenStudy (danjs):

the number you put is correct, but, i just see 8.3 * 25.4 to get it, not sure what you did

OpenStudy (fanduekisses):

Ohh thanks the way you explained it was very neat and clear! ^_^

OpenStudy (fanduekisses):

yes I skipped

OpenStudy (fanduekisses):

but the way you showed me is much proper

OpenStudy (danjs):

oh you have, 25.4 mm = 1 in, i dont have that remembered so yeah then 25.4^2 mm^2 = 1^2 in^2 645 mm^2 = 1 in^2

OpenStudy (danjs):

welcome, goodluck, yeah all ya need is the base unit to unit, and anything to change metric to US, and you are good, just less to remember